Knowing that mass of moon is M/81 iswhere M is the mass of earth, find the distance of the point from moon, where the net gravitational force due to earth and moon cancel each other. Given that distance between earth and moon is 60R, where Ris the radius of earth.

Grade:12th pass

2 Answers

Beenu Mathew
7 years ago
let there be a point mass of mass 1kg(m) between the earth and the moon.(and there is point where this mass doesnt experience any force.)
And this mass is at a distance of x from the moon.
and 60R-x from the earth.
so now lets dive in and get the fundamental equation
GM/81 * m/x2= GM*m/(60R-x)2
GM/81 * m/x2= GM*m/(60R-x)2
square root on both sides and skipping few steps
9x=60R-x
x=6R (from moon)
54R (from earth)
THANKS 
BEST OF LUCK!
PLS APPROVE
 
 
 
 
ritu
5 years ago
Let x be the distance between one body and the point where two fields cancel each other. And d be distance between two bodies .
Then,
Point where gravitational force is zero 
=√M(d)/√m+√M